minimum value of cos(cosx) is what , with explanation
 

Grade:11

2 Answers

Jayakrishna Natarajan
8 years ago
let y=cos(cosx)
dy/dx= sin(cosx)*sinx... dy/dx=0 @ x=2n(pi) or 2m+1(pi/2)
substituting this value for y..
y=cos1 or y=1
therefore range of y is [cos1,1]
jagdish singh singh
8 years ago
\hspace{-0.6 cm}$ Let $f(x)=\cos (\cos x)\;\forall x\in \mathbb{R}.$ Using $-1 \leq \cos x\leq 1$\\\\ So $ \cos 1 \leq \cos (\cos x)\leq 1.,$ Bcz $\cos x$ is an Decreasing function \\\\So above we have reverse Inequality and it is also an even function.\\\\So its value is same at two end points $\bf{i.e \; \cos(-1) = \cos (1)}$\\\\and Max. of $\cos$ is $\bf{1}.$ So $f(x)=\cos(\cos x)\in \left[\cos 1\;, 1\right].$